OpenStudy (anonymous):

The gas tank of a truck is a cylinder 5ft long with a diameter of 2ft . At the gas station, a pump pours gas at the rate of 3^ft3 per minute. How many minutes does it take to fill the empty tank with that pump? Use the value 3.14 for Pi , and round your answer to the nearest minute.

OpenStudy (anonymous):

V = pir^2h V= pi(1)^2(5) V = 5pi divide the volume by 3 to get the time

OpenStudy (anonymous):

thankyou